at maximum speed airplan travel 2460 miles against the wind in 6 hours flying with the wind the plan can travel the same distance in 5 hours
Answer by stanbon(75887)

at maximum speed an airplane can travel 2460 miles against the wind in 6 hours flying with the wind the plane can travel the same distance in 5 hours
----------------
Against the wind DATA:
distance = 2460 miles ; time = 6 hrs. ; rate = 2460/6 = 410 mph
----------------------------
With the wind DATA:
distance = 2460 miles ; time = 5 hrs. ; rate = 492 mph
----------------------------
Equations:
p + w = 492
p - w = 410
---------------------
Add:
2p = 902
p = 451 mph (speed of the plane in still air)
----------------
Substitute into p+w = 492 to solve for "w":
451+w = 492
w = 41 mph (speed of the wind)
================================================
Cheers,
Stan H.